# Interactions of the Pseudoscalar Meson Octet and the Baryon Decuplet in the Continuum and a Finite Volume

### v1

@@ II Resonances from UChPT
where $k$ and $k^{\prime}$ denote the momenta of the incoming and outgoing mesons, respectively. The constant $C_{ij}$ relies on the specific incoming and outgoing channels, and can be determined by utilizing the Lagrangian equation provided in Eq. (1) [^3] Note that several $C_{ij}$ are opposite to those in Ref. [15] due to the phase conventions in Eq. (2).. As mostly done in the literature, we neglect here the Born terms, that are formally also of leading order.
@@
For the case of $(S,I)=(0,1/2)$, there exist two channels, namely ($\pi\Delta,\ K\Sigma$), with thresholds (1370, 1880) MeV. Utilizing the coefficient matrix $C_{ij}$ where
@@
one can find two poles. One pole is located at $1387-77i$ MeV on RS (+)(-+)( - + ), directly linked to the physical region and resulting in a broad peak on the physical axis. In the RPP, the closest baryon to this pole is the $N(1520)$, situated approximately at $1510-55i$ MeV. We do not propose to match these states here as done in [15], as a better UChPT description based on Goldstone boson scattering of the baryon ground state octet can be obtained [31]. The second pole, found at $1883+12i$ MeV on RS (+)(+-)( + - ), is not directly connected to the physical region. Its influence on $|T_{K\Sigma\to K\Sigma}|$ is reflected as a peak-like threshold cusp exactly at the $K\Sigma$ threshold [32]. In the RPP, there is a nearby baryon $N(1875)$ located around $1900-80i$ MeV. The threshold cusp in our model provides a possible explanation for the experimental signal of the $N(1875)$. It is found in Refs. [33, 34] that there is no place for $N(1875)$ being a three quark state in the constituent quark model. A detailed study of $N(1875)$ from the $K\Sigma^{*}$ interaction, together with several coupled channels and possible triangle singularity, is presented in Ref. [35]. In Refs. [36, 37], the $N(1875)$ is identified as a $K\Sigma^{*}$ bound state in the one-boson-exchange model. A recent analysis [38] on the pion- and photon-induced production of $\pi N,\eta N,K\Lambda$ and $K\Sigma$ final states within the Jülich-Bonn model also indicates a dynamically generated state in $D_{13}$ partial wave, whose mass is close to $N(1875)$ but the total width ($\Gamma>600$ MeV) is much larger.
For the case of $(S,I)=(0,3/2)$, there are three channels, namely $\pi\Delta$, $\eta\Delta$ and $K\Sigma$, with corresponding thresholds at 1370, 1780, and 1880 MeV. Utilizing the matrix $C_{ij}$ where
@@
For the case of $(S,I)=(-1,0)$, there are two channels, namely, $\pi\Sigma$ and $K\Xi$, with corresponding thresholds of 1523 and 2029 MeV. By utilizing the matrix $C_{ij}$ given by
@@
For the case of $(S,I)=(-1,1)$, there are four channels, namely, $\pi\Sigma$, $\bar{K}\Delta$, $\eta\Sigma$ and $K\Xi$, with with respective thresholds at 1523, 1728, 1932 and 2029 MeV. By utilizing the matrix $C_{ij}$ given by
@@
For the case of $(S,I)=(-2,1/2)$, there are four channels, namely $\pi\Xi$, $\bar{K}\Sigma$, $\eta\Xi$ and $K\Omega$, respective thresholds at 1671, 1880, 2081, and 2168 MeV. By utilizing the matrix $C_{ij}$ given by
@@
For the case of $(S,I)=(-3,0)$, there are two channels, namely $\bar{K}\Xi$ and $\eta\Omega$, with corresponding thresholds at 2029 and 2220 MeV. By utilizing the matrix $C_{ij}$ given by
@@ IV.1 Flavor SU(3) symmetry
For the case of $(S,I)=(0,1/2)$, the relation between the isospin-symmetric particle basis ($\left|\pi\Delta\right\rangle,\ \left|K\Sigma\right\rangle$) and the SU(3) flavor basis ($\left|8\right\rangle,\ \left|27\right\rangle$) is given by
$\left(\begin{array}[]{c}|8\rangle\\ |27\rangle\end{array}\right)=U\left(\begin{array}[]{c}|\pi\Delta\rangle\\ |K\Sigma\rangle\end{array}\right),$ (57)
@@ IV.2 Fit results
To verify the consistency of the poles in the two scenarios, taking into account the uncertainties of the fitting, we conducted an error analysis using the $(S,I)=(0,1/2)$ sector as a case study. Since the energy levels were calculated using UChPT interaction without any errors, comparing the $\chi^{2}/\rm d.o.f.$ directly with 1 and interpreting the uncertainties of the fitted parameters would be inappropriate. It is worth mentioning that the energy level errors in Ref. [17] are on the order of 10 MeV, which is typically larger for baryons in lattice studies. Therefore, we assigned an error of 20 MeV to the energy levels of the $\Phi\bm{T}$ systems. In this particular case, for the $(S,I)=(0,1/2)$ sector, the two poles were found to be located at $1343\pm 11$ MeV on the RS (++) and at $1896^{+26}_{-19}-14^{+13}_{-9}i$ MeV on the RS (+)(-+)( - + ). The first pole in both cases exhibits agreement within the uncertainty, while the higher pole does not.
@@ V Summary
Next, we examine the interaction between $\Phi\bm{T}$ in a finite volume. To regularize the two point Green’s function, we introduce a form factor, which allows us to identify additional poles in these systems. These poles include several bound states that lie below the lowest threshold in each $(S,I)$ sector. Using UChPT as the underlying theory, we make predictions for the energy levels in finite volume with different lengths, denoted as $L$. These predicted energy levels are then fitted using an amplitude parameterized by the $K$-matrix, which satisfies flavor SU(3) symmetry. The utilization of SU(3) symmetry in the parameterization of the $K$-matrix helps reduce the number of parameters and improves the stability of the fit. The goodness of fit for the lowest four or five energy levels is satisfactory in each $(S,I)$ sector. By employing the parameters obtained from the best fits, we determine the pole positions of the $K$-matrix parameterized amplitude. We find that the $K$-matrix parametrization is able to accurately reproduce the poles of the underlying theory only when the pole is close to the physical region and its mass is not too far from the lowest relevant threshold. This suggests that when fitting the energy levels obtained from lattice study using the $K$-matrix parameterization, we should be mindful of the applicable energy region and be cautious when dealing with poles that are too high, as they may not be reliable.

### v2

@@ II Resonances from UChPT
where $k$ and $k^{\prime}$ denote the momenta of the incoming and outgoing mesons, respectively. The constant $C_{ij}$ relies on the specific incoming and outgoing channels, and can be determined by utilizing the Lagrangian equation provided in Eq. (1) [^3] Note that several $C_{ij}$ are opposite to those in Ref. [15] due to the phase conventions in Eq. (2).. As mostly done in the literature, we neglect here the Born terms, that are formally also of leading order. We refer to Refs. [26, 27] for some detailed discussions.
@@
For the case of $(S,I)=(0,1/2)$, there exist two channels, namely ($\pi\Delta,\ K\Sigma^{*}$), with thresholds (1370, 1880) MeV. Utilizing the coefficient matrix $C_{ij}$ where
@@
one can find two poles. One pole is located at $1387-77i$ MeV on RS (+)(-+)( - + ), directly linked to the physical region and resulting in a broad peak on the physical axis. In the RPP, the closest baryon to this pole is the $N(1520)$, situated approximately at $1510-55i$ MeV. We do not propose to match these states here as done in [15], as a better UChPT description based on Goldstone boson scattering of the baryon ground state octet can be obtained [26]. The second pole, found at $1883+12i$ MeV on RS (+)(+-)( + - ), is not directly connected to the physical region. Its influence on $|T_{K\Sigma^{*}\to K\Sigma^{*}}|$ is reflected as a peak-like threshold cusp exactly at the $K\Sigma^{*}$ threshold [33]. In the RPP, there is a nearby baryon $N(1875)$ located around $1900-80i$ MeV. The threshold cusp in our model provides a possible explanation for the experimental signal of the $N(1875)$. It is found in Refs. [34, 35] that there is no place for $N(1875)$ being a three quark state in the constituent quark model. A detailed study of $N(1875)$ from the $K\Sigma^{*}$ interaction, together with several coupled channels and possible triangle singularity, is presented in Ref. [36]. In Refs. [37, 38], the $N(1875)$ is identified as a $K\Sigma^{*}$ bound state in the one-boson-exchange model. A recent analysis [39] on the pion- and photon-induced production of $\pi N,\eta N,K\Lambda$ and $K\Sigma$ final states within the Jülich-Bonn model also indicates a dynamically generated state in $D_{13}$ partial wave, whose mass is close to $N(1875)$ but the total width ($\Gamma>600$ MeV) is much larger.
For the case of $(S,I)=(0,3/2)$, there are three channels, namely $\pi\Delta$, $\eta\Delta$ and $K\Sigma^{*}$, with corresponding thresholds at 1370, 1780, and 1880 MeV. Utilizing the matrix $C_{ij}$ where
@@
For the case of $(S,I)=(-1,0)$, there are two channels, namely, $\pi\Sigma^{*}$ and $K\Xi^{*}$, with corresponding thresholds of 1523 and 2029 MeV. By utilizing the matrix $C_{ij}$ given by
@@
For the case of $(S,I)=(-1,1)$, there are four channels, namely, $\pi\Sigma^{*}$, $\bar{K}\Delta$, $\eta\Sigma^{*}$ and $K\Xi^{*}$, with with respective thresholds at 1523, 1728, 1932 and 2029 MeV. By utilizing the matrix $C_{ij}$ given by
@@
For the case of $(S,I)=(-2,1/2)$, there are four channels, namely $\pi\Xi^{*}$, $\bar{K}\Sigma^{*}$, $\eta\Xi^{*}$ and $K\Omega$, respective thresholds at 1671, 1880, 2081, and 2168 MeV. By utilizing the matrix $C_{ij}$ given by
@@
For the case of $(S,I)=(-3,0)$, there are two channels, namely $\bar{K}\Xi^{*}$ and $\eta\Omega$, with corresponding thresholds at 2029 and 2220 MeV. By utilizing the matrix $C_{ij}$ given by
@@ IV.1 Flavor SU(3) symmetry
For the case of $(S,I)=(0,1/2)$, the relation between the isospin-symmetric particle basis ($\left|\pi\Delta\right\rangle,\ \left|K\Sigma^{*}\right\rangle$) and the SU(3) flavor basis ($\left|8\right\rangle,\ \left|27\right\rangle$) is given by
$\left(\begin{array}[]{c}|8\rangle\\ |27\rangle\end{array}\right)=U\left(\begin{array}[]{c}|\pi\Delta\rangle\\ |K\Sigma^{*}\rangle\end{array}\right),$ (57)
@@ IV.2 Fit results
To verify the consistency of the poles in the two scenarios, taking into account the uncertainties of the fit procedure, we conducted an error analysis using the $(S,I)=(0,1/2)$ sector as a case study. Since the energy levels were calculated using UChPT interaction without any errors, comparing the $\chi^{2}/\rm d.o.f.$ directly with 1 and interpreting the uncertainties of the fitted parameters would be inappropriate. It is worth mentioning that the energy level errors in Ref. [17] are on the order of 10 MeV, which is typically larger for baryons in lattice studies. Therefore, we assigned an error of 20 MeV to the energy levels of the $\Phi\bm{T}$ systems. For the particular case of the $(S,I)=(0,1/2)$ sector, the two poles were found to be located at $1343\pm 11$ MeV on the RS (++) and at $1896^{+26}_{-19}-14^{+13}_{-9}i$ MeV on the RS (+)(-+)( - + ). The first pole in both cases exhibits agreement within the uncertainty, while the higher pole does not, as shown in FIG. 7. We have used MINUIT [51, 52, 53] to fit the data and obtained the errors of the free parameters. According to the correlation matrix and errors, we can produce a number of samples. Then those sets of parameters within $1\sigma$ defined by theMahalanobis distance [54] are used to calculate the pole positions.
Figure 7: The distribution of the pole positions on RS (++)(++)( + + ) and (+)(-+)( - + ) for the systems with $(S,I)=(0,1/2)$ within 1 $\sigma$.
@@ V Summary
Next, we examine the interaction between $\Phi\bm{T}$ in a finite volume. To regularize the two point Green’s function, we introduce a form factor, which allows us to identify additional poles in these systems. These poles include several bound states that lie below the lowest threshold in each $(S,I)$ sector. Using UChPT as the underlying theory, we make predictions for the energy levels in finite volume with different lengths, denoted as $L$. These predicted energy levels are then fitted using an amplitude parameterized by the $K$-matrix, which satisfies flavor SU(3) symmetry. The utilization of SU(3) symmetry in the parameterization of the $K$-matrix helps reduce the number of parameters and improves the stability of the fit. The goodness of fit for the lowest four or five energy levels is satisfactory in each $(S,I)$ sector. By employing the parameters obtained from the best fits, we determine the pole positions of the $K$-matrix parameterized amplitude. We find that the $K$-matrix parametrization is able to accurately reproduce the poles of the underlying theory only when the pole is close to the physical region and its mass is not too far from the lowest relevant threshold. This suggests that when fitting the energy levels obtained from lattice study using the $K$-matrix parameterization, we should be mindful of the applicable energy region and be cautious when dealing with poles that are too high, as they may not be reliable. Note that if one wants to extend the range of applicability of the method discussed here, one has to include more terms (and poles) in the $K$-matrix parametrization, which also increases the numebr of fit parameters.